OpenStudy (chrisneedshelp):

Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum. 1/7=? + 3/14=?

OpenStudy (anonymous):

\[1/7+3/14=2/14+3/14=5/14\]

OpenStudy (chrisneedshelp):

1/7=? and 3/14=?

OpenStudy (chrisneedshelp):

Convert these unlike fractions to equivalent like fractions and add them

OpenStudy (anonymous):

1/7 = 2/14

OpenStudy (chrisneedshelp):

5/14 simplest form

OpenStudy (anonymous):

yes it is

OpenStudy (chrisneedshelp):

THANKS :)